# A system of NLS arising in optical material without Galilean symmetry

### v1

@@ 3 High frequency limit
[18, Theorem 2.1] Let $1\leq N\leq 3$, $\omega>0$. Then there exists at least one ground state $(u_{0},v_{0})$ to system (1.7), which is radially symmetric, $v_{0}$ is positive and $u_{0}$ is either positive or identically zero.
@@
Let $(\omega,c)\in\mathbb{R}\times\mathbb{R}^{N}$ satisfy $\omega>0$ and $c\neq 0$. Then
$\mu_{\omega,c}=|c|^{4-N}\mu_{\frac{\omega}{|c|^{2}},\frac{c}{|c|}}.$
@@
Let $\{(u_{n},v_{n})\}$ be a minimizing sequence for $\mu_{\omega,c}$, that is,
$N_{\omega,c}(u_{n},v_{n})=0,~{}~{}S_{\omega,c}\to\mu_{\omega,c},$
where $N_{\omega,c}$ and $S_{\omega,c}$ are defined in (2.2) and (1.10), respectively.
Let $(\tilde{u}_{n},\tilde{v}_{n})=\left(\frac{1}{|c|}u_{n}\left(\frac{x}{|c|}\right),\frac{1}{|c|}v_{n}\left(\frac{x}{|c|}\right)\right)$. Then
$N_{\frac{\omega}{|c|^{2}},\frac{c}{|c|}}(\tilde{u}_{n},\tilde{v}_{n})=|c|^{N-4}N_{\omega,c}(u_{n},v_{n})=0,$
$S_{\frac{\omega}{|c|^{2}},\frac{c}{|c|}}(\tilde{u}_{n},\tilde{v}_{n})=|c|^{N-4}S_{\omega,c}\to|c|^{N-4}\mu_{\omega,c}.$
Hence, we can obtain the desired result. ∎
@@
$(\tilde{u}_{\omega},\tilde{v}_{\omega})=\frac{1}{\omega^{\frac{1}{2}}}\left(u\left(\frac{x}{\omega^{\frac{1}{2}}}\right),v\left(\frac{x}{\omega^{\frac{1}{2}}}\right)\right).$ (3.1)
@@
$N_{1,\frac{c}{\sqrt{\omega}}}(\tilde{u}_{\omega},\tilde{v}_{\omega})=\frac{1}{\omega^{\frac{4-N}{2}}}N_{\omega,c}(u,v),~{}~{}S_{1,\frac{c}{\sqrt{\omega}}}(\tilde{u}_{\omega},\tilde{v}_{\omega})=\frac{1}{\omega^{\frac{4-N}{2}}}S_{\omega,c}(u,v).$
@@
$\mu_{\omega,c}=\omega^{\frac{4-N}{2}}\mu_{1,\frac{c}{\sqrt{\omega}}}.$
So, $(u,v)$ is a minimizer of $\mu_{\omega,c}$ is equivalent to the fact that $(\tilde{u},\tilde{v})$ is a minimizer of $\mu_{1,\frac{c}{\sqrt{\omega}}}$.
@@
Let $(u_{0},v_{0})$ be a radial minimizer of $\mu_{1,0}$ given by Lemma 3.1, then
$\frac{c}{\sqrt{\omega}}\cdot P(u_{0},v_{0})=0.$
@@
$N_{1,\frac{c}{\sqrt{\omega}}}(u_{0},v_{0})=0.$
@@
$\mu_{1,\frac{c}{\sqrt{\omega}}}\leq\mu_{1,0}.$ (3.2)
Hence, by the definition of $\tilde{u}_{\omega}$ and $\tilde{v}_{\omega}$ (see (3.1)), we can obtain that $(\tilde{u}_{\omega},\tilde{v}_{\omega})$ is uniformly bounded in $H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})$ as $|c|$ is bounded and $\omega\rightarrow+\infty$. ∎
In the next lemma we will give the relation between $\mu_{1,0}$ and $\mu_{1,\frac{c}{\sqrt{\omega}}}$.
@@
Let $1\leq N\leq 3$. Then $0\leq\mu_{1,0}-\mu_{1,\frac{c}{\sqrt{\omega}}}\rightarrow 0$ as $|c|$ is bounded and $\omega\rightarrow+\infty$.
@@
$\frac{N-2}{2}\int(|\nabla\tilde{u}_{w}|^{2}+|\nabla\tilde{v}_{w}|^{2})+\frac{N}{2}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)$
@@
$N_{1,\frac{c}{\sqrt{\omega}}}(\tilde{u}_{\omega},\tilde{v}_{\omega})=2Q_{1,\frac{c}{\sqrt{\omega}}}(\tilde{u}_{\omega},\tilde{v}_{\omega})-4D(\tilde{u}_{\omega},\tilde{v}_{\omega})=\omega^{\frac{N-4}{2}}N_{\omega,c}(u,v)=0.$ (3.4)
By the definition of $N_{1,\frac{c}{\sqrt{\omega}}}$ (see (2.2)), (3.4) is equivalent to
$\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})-4D(\tilde{u}_{\omega},\tilde{v}_{\omega})=0.$ (3.5)
Combining (3) and (3.5), we get
$(4-N)\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)=N\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+(N-2)\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega}).$ (3.6)
@@
Hence, by the Gagliardo-Nirenberg inequality (1.16), we
@@
$<$ $2Q_{1,\frac{c}{\sqrt{\omega}}}(\tilde{u}_{\omega},\tilde{v}_{\omega})=4D(\tilde{u}_{\omega},\tilde{v}_{\omega})$
@@
for $\omega$ is sufficiently large. Combining (3.6) and (3.7), then
@@
$\left(\tilde{u}^{\lambda}_{\omega},\tilde{v}^{\lambda}_{\omega}\right)=\left(\tilde{u}_{\omega}(x/\lambda),\tilde{v}_{\omega}(x/\lambda)\right),~{}~{}\lambda>0.$ (3.9)
@@
$N_{1,0}\left(\tilde{u}^{\lambda}_{\omega},\tilde{v}^{\lambda}_{\omega}\right)=\lambda^{N-2}\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\lambda^{N}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)-4\lambda^{N}D(\tilde{u}_{\omega},\tilde{v}_{\omega}),$
where $N_{1,0}$ is defined in (2.2) with $\omega=1$ and $c=0$.
From (3) and (3.6), we have
$N_{1,0}\left(\tilde{u}^{\lambda}_{\omega},\tilde{v}^{\lambda}_{\omega}\right)$
$=$ $\lambda^{N-2}\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)-\frac{\lambda^{N}}{4-N}\left(N\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+\frac{2c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})\right)$
@@
$N_{1,0}(\tilde{u}_{\omega},\tilde{v}_{\omega})=N_{1,\frac{c}{\sqrt{\omega}}}(\tilde{u}_{\omega},\tilde{v}_{\omega})-\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})=-\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega}).$
Case 1. If $\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})>0$, then $N_{1,0}(\tilde{u}_{\omega},\tilde{v}_{\omega})<0$. Obviously, using (3.8), if $|c|$ is bounded and $\omega\to+\infty$,
@@
$(N-2)\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})<\frac{2c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega}).$
@@
$N_{1,0}\left(\tilde{u}^{\lambda_{\omega}}_{\omega},\tilde{v}^{\lambda_{\omega}}_{\omega}\right)=0.$
Case 2. If $\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})<0$, then $N_{1,0}(\tilde{u}_{\omega},\tilde{v}_{\omega})>0$. Similarly,
@@
$N_{1,0}\left(\tilde{u}^{\lambda_{\omega}}_{\omega},\tilde{v}^{\lambda_{\omega}}_{\omega}\right)=0.$
@@
In fact, by assumption, $N\leq 3$, $|c|$ is bounded and $\omega\to+\infty$, and the estimates (3.7) and (3.8), we deduce that

### v2

@@ 3 High frequency limit
[17, Theorem 2.1] Let $1\leq N\leq 3$, $\omega>0$. Then there exists at least one ground state $(u_{0,\omega},v_{0,\omega})$ to system (1.7), which is radially symmetric, $v_{0,\omega}$ is positive and $u_{0,\omega}$ is either positive or identically zero.
@@
Let $(\omega,c)\in\mathbb{R}^{+}\times\mathbb{R}^{N}\backslash\{0\}$. Then
$\mu_{\omega,c,\alpha-3\gamma}=|\sigma|^{4-N}\mu_{\frac{\omega}{\sigma^{2}},\frac{c}{\sigma},\frac{\alpha-3\gamma}{\sigma^{2}}},~{}~{}\sigma>0.$
@@
Let $\{(u_{n},v_{n})\}$ be a minimizing sequence for $\mu_{\omega,c,\alpha-3\gamma}$, that is,
$N_{\omega,c,\alpha-3\gamma}(u_{n},v_{n})=0,~{}~{}S_{\omega,c,\alpha-3\gamma}\to\mu_{\omega,c,\alpha-3\gamma},$
where $N_{\omega,c,\alpha-3\gamma}$ and $S_{\omega,c,\alpha-3\gamma}$ are defined in (2.2) and (1), respectively.
Let $(\tilde{u}_{n},\tilde{v}_{n})=\left(\frac{1}{\sigma}u_{n}\left(\frac{x}{\sigma}\right),\frac{1}{\sigma}v_{n}\left(\frac{x}{\sigma}\right)\right)$, $\sigma>0$. Then
$N_{\omega,c,\alpha-3\gamma}(u_{n},v_{n})=$ $2\mathcal{Q}_{\omega,c,\alpha-3\gamma}(u_{n},v_{n})-4D(u_{n},v_{n})$
$=$ $2\Big{(}\frac{\sigma^{4-N}}{2}\int|\nabla\tilde{u}_{n}|^{2}+\frac{\sigma^{4-N}}{2}\int|\nabla\tilde{v}_{n}|^{2}+\frac{\omega\sigma^{2-N}}{2}\int|\tilde{u}_{n}|^{2}$
$+\frac{(3\gamma\omega+\alpha-3\gamma)\sigma^{2-N}}{2}\int|\tilde{v}_{n}|^{2}+\frac{\sigma^{3-N}}{2}c\cdot P(\tilde{u}_{n},\tilde{v}_{n})\Big{)}-4\sigma^{4-N}D(\tilde{u}_{n},\tilde{v}_{n})$
$=$ $2\sigma^{4-N}\Big{(}\frac{1}{2}\int|\nabla\tilde{u}_{n}|^{2}+\frac{1}{2}\int|\nabla\tilde{v}_{n}|^{2}+\frac{\omega\sigma^{-2}}{2}\int|\tilde{u}_{n}|^{2}$
$+\frac{(3\gamma\omega+\alpha-3\gamma)\sigma^{-2}}{2}\int|\tilde{v}_{n}|^{2}+\frac{\sigma^{-1}}{2}c\cdot P(\tilde{u}_{n},\tilde{v}_{n})\Big{)}-4\sigma^{4-N}D(\tilde{u}_{n},\tilde{v}_{n})$
$=$ $\sigma^{4-N}N_{\frac{\omega}{\sigma^{2}},\frac{c}{\sigma},\frac{\alpha-3\gamma}{\sigma^{2}}}(\tilde{u}_{n},\tilde{v}_{n}),$
and
$S_{\omega,c,\alpha-3\gamma}(u_{n},v_{n})=$ $\frac{\sigma^{4-N}}{2}\int|\nabla\tilde{u}_{n}|^{2}+\frac{\sigma^{4-N}}{2}\int|\nabla\tilde{v}_{n}|^{2}+\frac{\omega\sigma^{2-N}}{2}\int|\tilde{u}_{n}|^{2}$
$+\frac{(3\gamma\omega+\alpha-3\gamma)\sigma^{2-N}}{2}\int|\tilde{v}_{n}|^{2}+\frac{\sigma^{3-N}}{2}c\cdot P(\tilde{u},\tilde{v})-\sigma^{4-N}D(\tilde{u}_{n},\tilde{v}_{n})$
$=$ $\sigma^{4-N}S_{\frac{\omega}{\sigma^{2}},\frac{c}{\sigma},\frac{\alpha-3\gamma}{\sigma^{2}}}(\tilde{u}_{n},\tilde{v}_{n}).$
Hence, we have
$N_{\frac{\omega}{\sigma^{2}},\frac{c}{\sigma},\frac{\alpha-3\gamma}{\sigma^{2}}}(\tilde{u}_{n},\tilde{v}_{n})=|\sigma|^{N-4}N_{\omega,c,\alpha-3\gamma}(u_{n},v_{n})=0,$
$S_{\frac{\omega}{\sigma^{2}},\frac{c}{\sigma},\frac{\alpha-3\gamma}{\sigma^{2}}}(\tilde{u}_{n},\tilde{v}_{n})=|\sigma|^{N-4}S_{\omega,c,\alpha-3\gamma}(u_{n},v_{n})\to|c|^{N-4}\mu_{\omega,c,\alpha-3\gamma}.$
This means that we can obtain the desired result. ∎
@@
$(\tilde{u}_{\omega},\tilde{v}_{\omega})=\frac{1}{\omega^{\frac{1}{2}}}\left(u\left(\frac{x}{\omega^{\frac{1}{2}}}\right),v\left(\frac{x}{\omega^{\frac{1}{2}}}\right)\right),~{}~{}\omega>0.$ (3.1)
@@
$N_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}(\tilde{u}_{\omega},\tilde{v}_{\omega})=\frac{1}{\omega^{\frac{4-N}{2}}}N_{\omega,c,\alpha-3\gamma}(u,v),~{}~{}S_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}(\tilde{u}_{\omega},\tilde{v}_{\omega})=\frac{1}{\omega^{\frac{4-N}{2}}}S_{\omega,c,\alpha-3\gamma}(u,v).$
@@
$\mu_{\omega,c,\alpha-3\gamma}=\omega^{\frac{4-N}{2}}\mu_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}.$
So, $(u,v)$ is a minimizer of $\mu_{\omega,c,\alpha-3\gamma}$ is equivalent to the fact that $(\tilde{u},\tilde{v})$ is a minimizer of $\mu_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}$.
@@
Let $(u_{0,\omega},v_{0,\omega})$ be a radial minimizer of $\mu_{1,0,\frac{\alpha-3\gamma}{\omega}}$ given by Lemma 3.1, then
$\frac{c}{\sqrt{\omega}}\cdot P(u_{0,\omega},v_{0,\omega})=0.$
@@
$N_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}(u_{0,\omega},v_{0,\omega})=0.$
@@
$\mu_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}\leq\mu_{1,0,\frac{\alpha-3\gamma}{\omega}}.$ (3.2)
Next, we will consider the relation between $\mu_{1,0,\frac{\alpha-3\gamma}{\omega}}$ and $\mu_{1,0,0}$.
Let $(u_{0,0},v_{0,0})$ be a minimizer of $\mu_{1,0,0}$, where $\omega=1$ and $\alpha=3\gamma$ (see Lemma 3.1). From the Pohozaev identities (2.1) and (2.1), $(u_{0,0},v_{0,0})$ satisfies
$\frac{N-2}{2}\int(|\nabla u_{0,0}|^{2}+|\nabla v_{0,0}|^{2})+\frac{N}{2}\left(\|u_{0,0}\|_{L^{2}}^{2}+3\gamma\|v_{0,0}\|_{L^{2}}^{2}\right)=ND(u_{0,0},v_{0,0}).$ (3.3)
Since $(u_{0,0},v_{0,0})$ is the solution of (1.8) with $\omega=1$ and $\alpha=3\gamma$, then, from Lemma 3.2, we have
$N_{1,0,0}(u_{0,0},v_{0,0})=$ $0\displaystyle 2\mathcal{Q}_{1,0,0}(u_{0,0},v_{0,0})-4D(u_{0,0},v_{0,0})=$ 0 . (3.4)
By the definition of $N_{1,0,0}$ (see (2.2)), (3.4) is equivalent to
$\int\left(|\nabla u_{0,0}|^{2}+|\nabla v_{0,0}|^{2}\right)$ $+\left(\|u_{0,0}\|_{L^{2}}^{2}+3\gamma\|v_{0,0}\|_{L^{2}}^{2}\right)-4D(u_{0,0},v_{0,0})=0.$ (3.5)
Combining (3.3) and (3.5), we get
$(4-N)\int\left(|\nabla u_{0,0}|^{2}+|\nabla v_{0,0}|^{2}\right)=N\left(\|u_{0,0}\|_{L^{2}}^{2}+3\gamma\|v_{0,0}\|_{L^{2}}^{2}\right).$ (3.6)
Let
$\left(u_{0,0}^{\sigma}(x),v_{0,0}^{\sigma}(x)\right)=\left(u_{0,0}\left(\frac{x}{\sigma}\right),v_{0,0}\left(\frac{x}{\sigma}\right)\right),~{}~{}\sigma>0.$
Then
$N_{1,0,\frac{\alpha-3\gamma}{\omega}}\left(u_{0,0}^{\sigma},v_{0,0}^{\sigma}\right)=$ $\sigma^{N-2}\int\left(|\nabla u_{0,0}|^{2}+|\nabla v_{0,0}|^{2}\right)+\sigma^{N}\left(\|u_{0,0}\|_{L^{2}}^{2}+3\gamma\|v_{0,0}\|_{L^{2}}^{2}\right)$
$+\sigma^{N}\frac{\alpha-3\gamma}{\omega}\|v_{0,0}\|_{L^{2}}^{2}-4\sigma^{N}D(u_{0,0},v_{0,0}),$
where $N_{1,0,\frac{\alpha-3\gamma}{\omega}}$ is defined in (2.2).
From (3.3), (3.5) and (3.6), we have
$N_{1,0,\frac{\alpha-3\gamma}{\omega}}\left(u_{0,0}^{\sigma},v_{0,0}^{\sigma}\right)$
$=$ $\sigma^{N-2}\int\left(|\nabla u_{0,0}|^{2}+|\nabla v_{0,0}|^{2}\right)-\frac{N\sigma^{N}}{4-N}\left(\|u_{0,0}\|_{L^{2}}^{2}+3\gamma\|v_{0,0}\|_{L^{2}}^{2}\right)-\frac{\sigma^{N}}{4-N}\frac{N(\alpha-3\gamma)}{\omega}\|v_{0,0}\|_{L^{2}}^{2}$
$=$ $\frac{\sigma^{N-2}}{4-N}N\left(\|u_{0,0}\|_{L^{2}}^{2}+3\gamma\|v_{0,0}\|_{L^{2}}^{2}\right)-\frac{N\sigma^{N}}{4-N}(\left(\|u_{0,0}\|_{L^{2}}^{2}+3\gamma\|v_{0,0}\|_{L^{2}}^{2}+\frac{\alpha-3\gamma}{\omega}\|v_{0,0}\|_{L^{2}}^{2}\right).$
If $\alpha-3\gamma>0$, then there exists $\sigma\in(0,1)$ such that $N_{1,0,\frac{\alpha-3\gamma}{\omega}}\left(u_{0,0}^{\sigma},v_{0,0}^{\sigma}\right)=0$. If $\alpha-3\gamma<0$, then $\frac{\alpha-3\gamma}{\omega}\|v_{0,0}\|_{L^{2}}^{2}\to 0$ as $\omega\to\infty$. This implies that there exists $\sigma\in(1,2)$ such that $N_{1,0,\frac{\alpha-3\gamma}{\omega}}\left(u_{0,0}^{\sigma},v_{0,0}^{\sigma}\right)=0$ as $\omega\to\infty$. From the above argument, we can easily obtain that $\sigma\to 1$ as $\omega\to\infty$. In particular, if $\alpha-3\gamma=0$, obviously, $\sigma=1$ and $N_{1,0,\frac{\alpha-3\gamma}{\omega}}\left(u_{0,0}^{\sigma},v_{0,0}^{\sigma}\right)=0$.
Now, from the identity (3.4), we deduce that
$\sigma^{N-2}S_{1,0,\frac{\alpha-3\gamma}{\omega}}\left(u_{0,\omega}^{\sigma},v_{0,\omega}^{\sigma}\right)$
$=$ $\frac{1}{2}\int\left(|\nabla u_{0,0}|^{2}+|\nabla v_{0,0}|^{2}\right)+\frac{\sigma^{2}}{2}\left(\|u_{0,0}\|_{L^{2}}^{2}+3\gamma\|u_{0,0}\|_{L^{2}}^{2}\right)+\frac{\sigma^{2}(\alpha-3\gamma)}{\omega}\|v_{0,0}\|_{L^{2}}^{2}-\sigma^{2}D(u_{0,0},v_{0,0})$
$=$ $\frac{1}{2}\int\left(|\nabla u_{0,0}|^{2}+|\nabla v_{0,0}|^{2}\right)+\sigma^{2}\frac{1}{4}\left(\|u_{0,0}\|_{L^{2}}^{2}+3\gamma\|v_{0,0}\|_{L^{2}}^{2}\right)$
$-\frac{\sigma^{2}}{4}\int\left(|\nabla u_{0,0}|^{2}+|\nabla v_{0,0}|^{2}\right)+\frac{\sigma^{2}(\alpha-3\gamma)}{\omega}\|v_{0,0}\|_{L^{2}}^{2}$
$=$ $\left(\frac{1}{2}-\frac{\sigma^{2}}{4}\right)\int\left(|\nabla u_{0,0}|^{2}+|\nabla v_{0,0}|^{2}\right)+\sigma^{2}\frac{1}{4}\left(\|u_{0,0}\|_{L^{2}}^{2}+3\gamma\|v_{0,0}\|_{L^{2}}^{2}\right)+\frac{\sigma^{2}(\alpha-3\gamma)}{\omega}\|v_{0,0}\|_{L^{2}}^{2}$
$=$ $\frac{1}{2}\mathcal{Q}_{1,0,0}(u_{0,0},v_{0,0})+\frac{1}{4}(1-\sigma^{2})\int\left(|\nabla u_{0,0}|^{2}+|\nabla v_{0,0}|^{2}\right)+\frac{\sigma^{2}(\alpha-3\gamma)}{\omega}\|v_{0,0}\|_{L^{2}}^{2}$
$+\frac{1}{4}(\sigma^{2}-1)\left(\|u_{0,0}\|_{L^{2}}^{2}+3\gamma\|v_{0,0}\|_{L^{2}}^{2}\right)$
$=$ $\frac{1}{2}\mathcal{Q}_{1,0,0}(u_{0,0},v_{0,0})+\frac{\sigma^{2}(\alpha-3\gamma)}{\omega}\|v_{0,0}\|_{L^{2}}^{2}+o(1)$
$=$ $\mu_{1,0,0}+\frac{\sigma^{2}(\alpha-3\gamma)}{\omega}\|v_{0,0}\|_{L^{2}}^{2}+o(1),$
where we used the fact that $\mu_{1,0,0}=\frac{1}{2}\mathcal{Q}_{1,0,0}$ and $\mathcal{Q}_{1,0,0}$ is defined by (2.1). This means that
$\sigma^{N-2}\mu_{1,0,\frac{\alpha-3\gamma}{\omega}}=\mu_{1,0,0}+\frac{\sigma^{2}(\alpha-3\gamma)}{\omega}\|v_{0,0}\|_{L^{2}}^{2}+o(1).$
Then we deduce that
$\mu_{1,0,\frac{\alpha-3\gamma}{\omega}}\leq\mu_{1,0,0}+C,~{}~{}\text{as}~{}~{}\omega\to\infty,$ (3.7)
since $\sigma\to 1$ as $\omega\to\infty$. Combining (3.2) and (3.7), we can obtain that $(\tilde{u}_{\omega},\tilde{v}_{\omega})$ is uniformly bounded in $H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})$ as $|c|$ is bounded and $\omega\rightarrow+\infty$.
In the next lemma we will give the relation between $\mu_{1,0,0}$ and $\mu_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}$.
@@
Let $1\leq N\leq 3$. Then $\mu_{1,0,0}-\mu_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}\rightarrow 0$ as $|c|$ is bounded and $\omega\rightarrow+\infty$.
@@
$\frac{N-2}{2}\int(|\nabla\tilde{u}_{w}|^{2}+|\nabla\tilde{v}_{w}|^{2})+\frac{N}{2}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+\frac{N}{2}\frac{\alpha-3\gamma}{\omega}\|\tilde{v}_{\omega}\|_{L^{2}}^{2}$
@@
$N_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}(\tilde{u}_{\omega},\tilde{v}_{\omega})=2\mathcal{Q}_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}(\tilde{u}_{\omega},\tilde{v}_{\omega})-4D(\tilde{u}_{\omega},\tilde{v}_{\omega})=\omega^{\frac{N-4}{2}}N_{\omega,c,\alpha-3\gamma}(u,v)=0.$ (3.9)
By the definition of $N_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}$ (see (2.2)), (3.9) is equivalent to
$\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)$ $+\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+\frac{\alpha-3\gamma}{\omega}\|\tilde{v}_{\omega}\|_{L^{2}}^{2}$
$+\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})-4D(\tilde{u}_{\omega},\tilde{v}_{\omega})=0.$ (3.10)
Combining (3) and (3), we get
$(4-N)\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)$
$=$ $N\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+\frac{N(\alpha-3\gamma)}{\omega}\|\tilde{v}_{\omega}\|_{L^{2}}^{2}+(N-2)\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega}).$ (3.11)
@@
Hence, by the Gagliardo-Nirenberg inequality (1.16), we get
@@
$<$ $2Q_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}(\tilde{u}_{\omega},\tilde{v}_{\omega})=4D(\tilde{u}_{\omega},\tilde{v}_{\omega})$
$\leq$ $4C_{opt}^{(1)}K(\tilde{u}_{\omega},\tilde{v}_{\omega})^{\frac{N}{2}}M_{3\gamma}(\tilde{u}_{\omega},\tilde{v}_{\omega})^{2-\frac{N}{2}}$
@@
for $\omega$ is sufficiently large. Combining (3) and (3.12), then
@@
Also, we have
$\frac{N(\alpha-3\gamma)}{\omega}\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\to 0,~{}~{}\text{as}~{}~{}\omega\to\infty.$ (3.14)
@@
$\left(\tilde{u}^{\lambda}_{\omega}(x),\tilde{v}^{\lambda}_{\omega}(x)\right)=\left(\tilde{u}_{\omega}(x/\lambda),\tilde{v}_{\omega}(x/\lambda)\right),~{}~{}\lambda>0.$ (3.15)
@@
$N_{1,0,0}\left(\tilde{u}^{\lambda}_{\omega},\tilde{v}^{\lambda}_{\omega}\right)=$ $\lambda^{N-2}\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)+\lambda^{N}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)-4\lambda^{N}D(\tilde{u}_{\omega},\tilde{v}_{\omega}),$
where $N_{1,0,0}$ is defined in (2.2).
From (3), (3) and (3), we have
$N_{1,0,0}\left(\tilde{u}^{\lambda}_{\omega},\tilde{v}^{\lambda}_{\omega}\right)$
$=$ $\lambda^{N-2}\int\left(|\nabla\tilde{u}_{\omega}|^{2}+|\nabla\tilde{v}_{\omega}|^{2}\right)-\frac{N\lambda^{N}}{4-N}\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}+\frac{2c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})\right)$
@@
$N_{1,0,0}(\tilde{u}_{\omega},\tilde{v}_{\omega})=$ $N_{1,\frac{c}{\sqrt{\omega}},\frac{\alpha-3\gamma}{\omega}}(\tilde{u}_{\omega},\tilde{v}_{\omega})-\frac{\alpha-3\gamma}{\omega}\|\tilde{v}_{\omega}\|_{L^{2}}^{2}-\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})$
$=$ $-\frac{\alpha-3\gamma}{\omega}\|\tilde{v}_{\omega}\|_{L^{2}}^{2}-\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega}).$
Case 1. If $\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})>0$ and $\alpha-3\gamma>0$, then $N_{1,0,\frac{\alpha-3\gamma}{\omega}}(\tilde{u}_{\omega},\tilde{v}_{\omega})<0$. Obviously, using (3.13), if $|c|$ is bounded and $\omega\to+\infty$,
@@
$(N-2)\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})<\frac{2c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega}),~{}~{}1\leq N\leq 3.$
@@
$N_{1,0,0}\left(\tilde{u}^{\lambda_{\omega}}_{\omega},\tilde{v}^{\lambda_{\omega}}_{\omega}\right)=0.$
Case 2. If $\frac{c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})<0$ and $\alpha-3\gamma<0$, then $N_{1,0,0}(\tilde{u}_{\omega},\tilde{v}_{\omega})>0$. Similarly,
@@
$N\left(\|\tilde{u}_{\omega}\|_{L^{2}}^{2}+3\gamma\|\tilde{v}_{\omega}\|_{L^{2}}^{2}\right)+\frac{2c}{\sqrt{\omega}}\cdot P(\tilde{u}_{\omega},\tilde{v}_{\omega})>0.$
@@
$N_{1,0,0}\left(\tilde{u}^{\lambda_{\omega}}_{\omega},\tilde{v}^{\lambda_{\omega}}_{\omega}\right)=0.$
@@
In fact, by assumption, $1\leq N\leq 3$, $|c|$ is bounded and $\omega\to+\infty$, and the estimates (3.12) and (3.13), we deduce that